OpenStudy (anonymous):

what is the zero(s) for the function f(x) = (x – 15)(x + 1)(x – 10) ?

OpenStudy (anonymous):

is that it? i thought there was more work involved

OpenStudy (anonymous):

15, -1, and 10

OpenStudy (anonymous):

u could ask others :)

OpenStudy (anonymous):

zero times the rest of the stuff will be zero. set each part equal to zero and solve. That's how we came up with those numbers.

OpenStudy (anonymous):

x-15=0 therefore x=15 x+1=0 therefore x=-1 x-10=0 therefore x=10